# Quantum Entanglement Autodistillation in Baryon Pair Decays

### v1

@@ 6 Mechanism Behind Entanglement Increase in Particle Decays
In this section, we discuss the underlying mechanism responsible for the spontaneous increase in entanglement observed after the decay of certain systems. For the decay of a spin-1/2 mother particle into a spin-1/2 daughter particle and a spin-0 meson, the relationship between the polarization vectors of the mother and daughter particles can be characterized by the decay parameters $\alpha_{D}$, $\beta_{D}$, and $\gamma_{D}$, as given in Ref. [42]:
@@
where $\mathbf{P}_{M}$ and $\mathbf{P}_{D}$ denote the polarization vectors of the mother and daughter particles in their respective rest frames, and $\hat{\mathbf{n}}$ is the unit vector in the direction of the daughter particle, defined in the rest frame of the mother particle.
Consider a mother particle with spin-up polarization, i.e., $\mathbf{P}_{M}=\hat{\mathbf{z}}$. After decay, the polarization of the daughter particle in its helicity frame is given by the vector $(\beta_{D},\gamma_{D},\alpha_{D}+\hat{\mathbf{z}}\cdot\hat{\mathbf{n}})$. Similarly, if the mother particle is spin-down ($\mathbf{P}_{M}=-\hat{\mathbf{z}}$), the daughter particle’s polarization becomes $(-\beta_{D},-\gamma_{D},\alpha_{D}-\hat{\mathbf{z}}\cdot\hat{\mathbf{n}})$.
@@
$(\beta_{D},\gamma_{D},\alpha_{D}+\hat{z}\cdot\hat{n})\quad\text{and}\quad(-\beta_{D},-\gamma_{D},\alpha_{D}-\hat{z}\cdot\hat{n}),$
@@
$\tan\phi_{1}$ $=\frac{\gamma_{D}}{\beta_{D}}=\frac{\cos\phi_{D}}{\sin\phi_{D}},$
$\tan\phi_{2}$ $=\frac{-\gamma_{D}}{-\beta_{D}}=\frac{\cos\phi_{D}}{\sin\phi_{D}},$
$\cos\theta_{1}$ $=\frac{\alpha_{D}+\hat{z}\cdot\hat{n}}{1+2\alpha_{D}\hat{z}\cdot\hat{n}+(\hat{z}\cdot\hat{n})^{2}},$
$\cos\theta_{2}$ $=\frac{\alpha_{D}-\hat{z}\cdot\hat{n}}{1-2\alpha_{D}\hat{z}\cdot\hat{n}+(\hat{z}\cdot\hat{n})^{2}}.$
@@ Acknowledgements
We are grateful to Prof. Zhaofeng Kang, Jianwei Cui, and Ruilin Zhu for their valuable discussions and insightful suggestions, which greatly contributed to the development of this work. Special thanks go to Guomengchao Bian for many enjoyable mathematical conversations that helped broaden our perspective during the course of this research. This work is supported by the Natural Science Foundation of China under Grant No. 12375086 and No.12005070, and Hubei Provincial Natural Science Foundation of China under Grant No.2025AFB557.

### v2

@@ 6 Mechanism Behind Entanglement Increase in Particle Decays
In this section, we discuss the underlying mechanism responsible for the spontaneous increase in entanglement observed after the decay of certain systems. For the decay of a spin-1/2 mother particle into a spin-1/2 daughter particle and a spin-0 meson, the relationship between the polarization vectors of the mother and daughter particles can be characterized by the decay parameters $\alpha_{D}$, $\beta_{D}$, and $\gamma_{D}$, as given in Ref. [46, p. 1001]:
@@
where $\mathbf{P}_{M}$ and $\mathbf{P}_{D}$ denote the polarization vectors of the mother and daughter particles in their respective rest frames, and $\hat{\mathbf{n}}$ is the unit vector in the direction of the daughter particle, defined in the rest frame of the mother particle. This expression was mentioned in Lee’s seminal work [44], although no explicit derivation was provided there. It can be derived using the density matrix formalism applied to the decay process [42], which in turn is based on the helicity formalism developed by Jacob and Wick [49]. A detailed derivation of the polarization formula can be found in the Appendix.
Consider a mother particle with spin-up polarization, i.e., $\mathbf{P}_{M}=\hat{\mathbf{z}}$. After decay, the polarization of the daughter particle in its helicity frame is given by the vector $(\beta_{D}\sin{\theta},\,\gamma_{D}\sin{\theta},\,\alpha_{D}+\cos{\theta})$, where $\cos{\theta}=\hat{\mathbf{z}}\cdot\hat{\mathbf{n}}$. Similarly, if the mother particle is spin-down ($\mathbf{P}_{M}=-\hat{\mathbf{z}}$), the daughter particle’s polarization becomes $(-\beta_{D}\sin{\theta},\,-\gamma_{D}\sin{\theta},\,\alpha_{D}-\cos{\theta})$.
@@
$(\beta_{D}\sin{\theta},\,\gamma_{D}\sin{\theta},\,\alpha_{D}+\cos{\theta})\quad\text{and}\quad(-\beta_{D}\sin{\theta},\,-\gamma_{D}\sin{\theta},\,\alpha_{D}-\cos{\theta}),$
@@
$\tan\phi_{1}$ $=\frac{\gamma_{D}\sin\theta}{\beta_{D}\sin\theta}=\frac{\cos\phi_{D}}{\sin\phi_{D}},$
$\tan\phi_{2}$ $=\frac{-\gamma_{D}\sin\theta}{-\beta_{D}\sin\theta}=\frac{\cos\phi_{D}}{\sin\phi_{D}},$
$\cos\theta_{1}$ $=\frac{\alpha_{D}+\cos\theta}{1+\alpha_{D}\cos\theta},$
$\cos\theta_{2}$ $=\frac{\alpha_{D}-\cos\theta}{1-\alpha_{D}\cos\theta}.$
@@ Acknowledgements
We are grateful to Zhaofeng Kang, Jianwei Cui, and Ruilin Zhu for their valuable discussions and insightful suggestions, which greatly contributed to the development of this work. Special thanks go to Guomengchao Bian for many enjoyable mathematical conversations that helped broaden our perspective during the course of this research. This work is supported by the Natural Science Foundation of China under Grant No. 12375086 and No.12005070, and Hubei Provincial Natural Science Foundation of China under Grant No.2025AFB557.
Appendix
Derivation of the Polarization Formula
The density matrix of a spin-up mother particle can be expressed as
$\rho_{M}=|\uparrow\rangle\langle\uparrow|=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=\frac{1}{2}\sigma_{0}+\frac{1}{2}\sigma_{3}.$
With the density matrix formalism mentioned in [42], the density matrix of the spin of the final particles in their own helicity frame can be expressed as
$\rho_{D}$ $=\frac{1}{2}\sum_{\mu=0}^{3}a_{0\mu}\sigma_{\mu}+\frac{1}{2}\sum_{\mu=0}^{3}a_{3\mu}\sigma_{\mu}$
$=\frac{1}{2}[(1+\alpha_{D}\cos{\theta})\sigma_{0}-\gamma_{D}\sin{\theta}\sigma_{1}+\beta_{D}\sin{\theta}\sigma_{2}+(\alpha_{D}+\cos{\theta})\sigma_{3}].$
As we mentioned in Section 5, this unnormalized density matrix contains two parts of information, the spin configuration and the angular distribution of the daughter particle. The spin configuration of the daughter particle can be obtained by normalizing the density matrix
$\rho_{D}^{\prime}=\frac{\rho_{D}}{Tr(\rho_{D})}=\frac{1}{2}(\sigma_{0}+\vec{r}\cdot\vec{\sigma}),$
where
$\vec{r}=\frac{1}{1+\alpha_{D}\cos{\theta}}(-\gamma_{D}\sin{\theta},\beta_{D}\sin{\theta},\alpha_{D}+\cos{\theta}).$
The density matrix $\rho_{D}^{\prime}$ represents a pure state, and the vector $\vec{r}$ denotes the polarization direction of the daughter particle [50, p. 105]. This result is consistent with the expression given by Lee, in which the daughter polarization direction is $(\beta_{D}\sin{\theta},\ \gamma_{D}\sin{\theta},\ \alpha_{D}+\cos{\theta})$, up to a rotation of $\pi/2$ around the $\hat{z}-$axis between the coordinate systems used. This difference arises solely from the choice of coordinate convention and does not affect any physical predictions.
The polarization formula can also be derived by Feynman’s formalism [51, pp. 223–225]. With this method, we can derive the transition rate as
$R(\hat{\omega}_{i},\hat{\omega}_{f},\hat{\mathbf{n}})=1+\gamma_{D}\,\hat{\omega}_{f}\cdot\hat{\omega}_{i}+(1-\gamma_{D})(\hat{\omega}_{f}\cdot\hat{\mathbf{n}})(\hat{\omega}_{i}\cdot\hat{\mathbf{n}})+\alpha_{D}\left(\hat{\omega}_{f}\cdot\hat{\mathbf{n}}+\hat{\omega}_{i}\cdot\hat{\mathbf{n}}\right)+\beta_{D}\,\hat{\mathbf{n}}\cdot(\hat{\omega}_{f}\times\hat{\omega}_{i}),$
Here, $\hat{\mathbf{n}}$ is the unit vector in the direction of the final baryon’s momentum, and $\hat{\omega}_{i}$, $\hat{\omega}_{f}$ are the spin direction unit vectors of the initial and final baryons. By choosing
$\hat{\omega}_{f}=x_{f}\,\hat{\omega}_{i}\times\hat{\mathbf{n}}+y_{f}\,\hat{\mathbf{n}}\times(\hat{\omega}_{i}\times\hat{\mathbf{n}})+z_{f}\,\hat{\mathbf{n}}$,
the transition rate becomes
$R(\hat{\omega}_{i},\hat{\omega}_{f},\hat{\mathbf{n}})=1+\alpha_{D}\cos\theta+(\cos{\theta}+\alpha_{D})z_{f}+\beta_{D}x_{f}\sin^{2}{\theta}+\gamma_{D}y_{f}\sin^{2}{\theta},$ (12)
where
$\cos\theta=\hat{\omega}_{i}\cdot\hat{\mathbf{n}}$.
Now, if we choose
$(x_{f},y_{f},z_{f})=-\frac{1}{1+\alpha_{D}\cos\theta}(\beta_{D},\gamma_{D},\alpha_{D}+\cos\theta),$
then
$R(\hat{\omega_{i}},\hat{\omega_{f}},\hat{\mathbf{n}})=0.$
Thus, the opposite polarization direction, corresponding to an orthogonal spin state, will be the polarization direction of the daughter baryon:
$\hat{\omega}_{f}=\frac{(\alpha_{D}+\cos\theta)\hat{\mathbf{n}}+\beta_{D}\hat{\omega}_{i}\times\hat{\mathbf{n}}+\gamma_{D}\hat{\mathbf{n}}\times(\hat{\omega}_{i}\times\hat{\mathbf{n}})}{1+\alpha_{D}\cos\theta}.$
Alternatively, this result can also be derived by finding the extreme value of Eq. 12, subject to the constraint
$x_{f}^{2}\sin^{2}\theta+y_{f}^{2}\sin^{2}\theta+z_{f}^{2}=1.$